OpenStudy (anonymous):

Prove that if a and b are odd integers, then 4/(a^2+b^2) (proof by contradiction)

OpenStudy (anonymous):



OpenStudy (anonymous):

Im actually just looking for the initial negation of the statement.

OpenStudy (anonymous):

if a and b are even, then (a^2+b^2) is divisible by 4?

OpenStudy (espex):

Yes.

OpenStudy (espex):

But you need to assume the contrary of your initial state.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

So the negation (contrary) of the statement if a and b are even, then (a^2+b^2) is divisible by 4?

OpenStudy (espex):

You start with, if a & b are odd, then 4 is divisible by \(a^2+b^2\), was there more to that statement?

OpenStudy (anonymous):

*is not if***

OpenStudy (anonymous):

Original statement that I am proving is as follows: If a & b are off, then a^2+b^2 is NOT divisible by 4.

OpenStudy (anonymous):

I am thinking that the negation of this statement is as follows: If a & b are even, then a^2+b^2 IS divisible by 4.

OpenStudy (anonymous):

**odd, not off**

OpenStudy (anonymous):

 this notation means (a^2+b^2) is NOT divisible by 4

OpenStudy (anonymous):

a^2+b^2 mod 4

OpenStudy (espex):

Proofs are not my strong suit, I think @amistre64 might be better suited to help you. That said, if your original statement was that they are NOT divisible, then you must assume the contrary, that they ARE.

OpenStudy (anonymous):

Okay but do not change the even or off part of the statement?

OpenStudy (anonymous):

****Odd

OpenStudy (espex):

No, because you are trying to prove the contrary of the original condition, which was that odd numbers were not divisible by 4.

OpenStudy (anonymous):

okay I will work with that. Thanks

OpenStudy (amistre64):

if a and b are odd integers, then 4/(a^2+b^2) ... is that spose to be 4 divides (...) now contradict it if a and b are odd integers, then 4\(\not |\) (a^2+b^2)

OpenStudy (amistre64):

let a = 2n+1, and b=2k-1 be odd integers for any n,k in \(Z\)

OpenStudy (amistre64):

that -1 on the end of b is not going to be product so lets get it to +1 instead \[a^2+b^2 =(2n+1)^2+(2k+1)^2\] \[... =(4n^2+4n+1)+(4k^2-4k+1)\] \[... =4n^2+4n+1+4k^2+4k+1\] \[... =4(n^2+n+k^2+k)+2\]

OpenStudy (amistre64):

lol, forgot to edit it on the second line :) +4k it is

OpenStudy (amistre64):

i wonder if completing the squares onm n and k would be useful

OpenStudy (amistre64):

\[... =4(n^2+n+k^2+k)+2\] \[... =4(n^2+n+\frac14-\frac14+k^2+k+\frac14-\frac14)+2\] \[... =4((n+\frac12)^2+(k+\frac12)^2-\frac12))+2\] \[... =4((n+\frac12)^2+(k+\frac12)^2)-2+2\] \[... =4((n+\frac12)^2+(k+\frac12)^2)\] it should be apparent now that 4 divides it which contradicts the setup

OpenStudy (amistre64):

let m = that sum of squares jargon \[a^2 + b^2 = 4m\] since 4 divides the right side, 4 also divides the left side